# Emulating Full Client Participation: A Long-Term Client Selection Strategy for Federated Learning

### v1

@@ Abstract
Client selection significantly affects the system convergence efficiency and is a crucial problem in federated learning. Existing methods often select clients by evaluating each round individually and overlook the necessity for long-term optimization, resulting in suboptimal performance and potential fairness issues. In this study, we propose a novel client selection strategy designed to emulate the performance achieved with full client participation. In a single round, we select clients by minimizing the gradient-space estimation error between the client subset and the full client set. In multi-round selection, we introduce a novel individual fairness constraint, which ensures that clients with similar data distributions have similar frequencies of being selected. This constraint guides the client selection process from a long-term perspective. We employ Lyapunov optimization and submodular functions to efficiently identify the optimal subset of clients, and provide a theoretical analysis of the convergence ability. Experiments demonstrate that the proposed strategy significantly improves both accuracy and fairness compared to previous methods while also exhibiting efficiency by incurring minimal time overhead.
@@
Federated Learning, Client Selection, Individual Fairness, Coreset Selection, Lyapunov Optimization
@@ I Introduction
Federated learning (FL) facilitates collaborative model training without the necessity of sharing local data [1, 2], and is widely used in various domains [3, 4, 5]. In FL, model parameters or gradient updates are frequently exchanged between the central server and participating clients, leading to substantial communication overhead. To address the challenge posed by the limited bandwidth, a common approach is to select a subset of clients for local training, which reduces the communication costs compared with involving all clients [6, 7]. However, in heterogeneous scenarios, different selection strategies lead to distinct optimization trajectories, and an improper client selection strategy can cause the optimization trajectory to deviate seriously from the optimal one, significantly hindering the model performance. Therefore, a critical challenge is to develop an effective client selection strategy that reduces the communication overhead while ensuring model convergence.
In most cases, the global models generated by full client participation are considered the gold standard or benchmark. This is because they represent models trained on the entire dataset available across all clients, offering a comprehensive reference point for evaluating model performance. Building upon this understanding, we propose an intuitive strategy where clients are selected in a manner that ensures the optimization trajectory closely resembles that achieved with full client participation over the long term. We refer to this strategy as the long-term client selection strategy. This strategy yields two key advantages. First, by emulating the optimization trajectory of full client participation, it aims to achieve comparable performance. Moreover, it reduces communication overhead by involving only a subset of clients in the process.
Our Goal. In this work, we investigate the three questions: 1) what principles should be followed to implement this strategy; 2) how to mathematically formulate this strategy and provide a fast solution; and 3) is this strategy more effective than existing methods? To answer the first question, we propose two crucial principles to be followed.
Principle I: In each round, the data distribution generated by the selected subset of clients should closely resemble the data distribution of the full client participation. For example, in Fig. 1 (a), clients collaboratively train an image classification model, where each client has images of only one kind of animal, and three clusters exist in their data distributions. Due to limited bandwidth, the server is required only to select three clients in each round. In this case, as shown in Fig. 1 (b), the ideal strategy is to select one client from each cluster instead of three clients from the same cluster. This ensures that the data distribution resulting from the subset of selected clients serves as a reliable estimation of that from the full set of clients. However, in real-world applications, the clustering pattern in data distributions may be unknown or non-existent. Therefore, the key challenge here is to determine the subset of clients that best approximates the full client set without imposing extensive assumptions on their data distributions.
Figure 1: An example of the long-term client selection strategy.
However, if only Principle I is followed and client selection is performed by evaluating each round individually, there is a risk that the same subset of clients may be consistently selected across multiple rounds, while others may have minimal opportunities for selection. This phenomenon is indeed observed in our experimental results. In such cases, the system may accurately predict outcomes for frequently selected clients but makes inaccurate or biased predictions for those rarely selected, resulting in degradation of the global model’s performance and ineffective convergence. Furthermore, these seldom-selected clients may perceive unfair treatment and lose motivation to participate, ultimately endangering the sustainability of the federated system.
To mitigate this problem, we propose Principle II, which guides client selection across multiple rounds from a long-term perspective. Principle II: In multi-round selection, clients with similar data distributions should have similar frequencies of being selected. For example, as illustrated in Fig. 1, clients 1, 2, and 3 should share similar selection frequencies, as should clients 4 and 5, while the frequencies for clients 1 and 4 may differ. This principle aligns with the concept of individual fairness (IF) in central machine learning [8], which posits that similar instances or individuals in a dataset should yield similar predictions or outcomes. In our work, we refer to this principle as the individual fairness constraint. Here, the key challenge is to quantify the similarity between clients in terms of data distribution and selection frequency, respectively, and further formulate individual fairness mathematically.
This proposed individual fairness offers two key benefits. First, as illustrated in Fig. 1 (c), it promotes the diversity of selected clients across multiple rounds and reduces model prediction bias. Furthermore, existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients. However, this uniform approach may compromise selection performance, as it disregards the heterogeneous data distribution among clients. In contrast, our proposed fairness constraint specifically targets clients with similar data distributions. Thus, it allows for varying selection frequencies among clients with heterogeneous data distributions, such as clients 1 and 4 in Fig. 1. This flexibility enables the individual constraint to accommodate the data heterogeneity without compromising selection performance in single rounds.
Limitation of Prior Works. To address the client selection problem, a common approach is the random selection strategy [1]. Recent works have proposed various improvements, including contribution-based [12, 1, 13, 14, 15], loss-based [16, 17, 18], and cluster-based methods [19, 20]. However, these methods typically evaluate each round individually and fail to select clients from a long-term perspective, potentially leading to model biases and fairness issues. Additionally, the clustering pattern in data distributions may be unknown or non-existent, making the application of cluster-based methods challenging. Furthermore, fair federated learning has discussed collaborative fairness [13, 15, 21], group fairness [22, 22] and performance fairness [23, 24], where performance fairness aligns most closely with our objectives. Existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients. However, it disregards the heterogeneous data distribution among clients and requires their selection probabilities to be uniformly similar. This may hinder the identification of the optimal subset of clients that best approximates the full set of clients.
In this study, we propose a novel long-term client selection strategy in federated learning (LongFed), designed to emulate the model performance achieved with full client participation. In a single round, we perform client selection by minimizing the estimation error between a representative subset of clients and the full set of clients. In multi-round selection, we propose a novel individual fairness constraint to promote the diversity of selected clients without compromising selection performance in single rounds. Subsequently, we formulate the client selection problem as a time-averaged optimization function and simplify it using Lyapunov optimization from control theory. In addition, leveraging the submodular nature of the optimization function, we propose a fast greedy algorithm to find the optimal subset of clients in each round. The convergence of the proposed strategy is also analyzed. We conduct extensive experiments to validate the proposed LongFed strategy. First, the proposed strategy effectively guides the system to converge along a trajectory similar to that of full client participation. Second, it outperforms prior works in terms of both accuracy and fairness. Last, the proposed strategy introduces minimal time overhead compared to the random selection strategy. Our contributions are as follows.
  • We propose a novel long-term client selection strategy designed to guide the system to converge along a trajectory similar to that of full client participation. This strategy does not require extensive assumptions on the data distribution.
@@
  • We identify and address an issue that has been rarely discussed in previous works, i.e., the frequency of selecting clients with similar distributions. We propose an individual fairness criterion to mathematically formulate this issue and employ Lyapunov optimization to solve it.
@@ II-A Client Selection in Federated Learning
However, most of these methods select clients by evaluating each round individually, thus failing to optimize a series of client subsets from a long-term perspective. These limitations could cause potential model biases and fairness issues, ultimately resulting in lower convergence efficiency. Furthermore, if the clustering pattern among clients is known, the long-term client selection problem becomes straightforward: simply selecting one client randomly from each cluster in every communication round [20]. However, in practical scenarios, the clustering pattern is usually unknown or, in some cases, does not exist. This limitation renders cluster-based methods ineffective for solving the long-term client selection problem.
@@ II-B Fairness in Federated Learning
Performance fairness refers that the model should produce similar performance across all clients, where the uniformity of performance is commonly used as the metric for evaluation [30, 31]. Various approaches have emerged to achieve performance fairness. For example, in [24], fairness is defined by assigning uniform weight to each client and introduced as an additional constraint in the optimization function. Besides, [23] identifies a trade-off between model robustness and fairness, and proposes a personalized framework to inherently achieve both fairness and robustness benefits. Moreover, several works study the fairness issue by considering that the probability of being selected is similar for all clients [9, 10, 11].
However, the three types of fairness described above cannot effectively address the fairness issue outlined in Principle II. While collaborative fairness and group fairness focus on different aspects compared to our work, performance fairness aligns most closely with our objectives. In these works [9, 10, 11], the uniform sampling method ensures a similar probability of being selected for all clients. However, we argue that it is unsuitable for achieving the long-term client selection strategy. This is because the uniform sampling method disregards the heterogeneous data distribution among clients. Consequently, it may hinder the identification of the optimal subset of clients that best approximates the data distribution of full client participation.
@@ III The Proposed Optimization Function
In this section, we begin with an introduction to the federated learning system. Next, we illustrate the client selection in a single round and across multiple rounds, followed by the proposed long-term optimization function. Last, we apply Lyapunov optimization to transform and simplify the optimization function.
@@ III-A Preliminary of Federated Learning
In our work, we consider that the federated learning system consists of a central server and a set of clients denoted as =$\mathbb{N}=\{1,\cdots,N\}$. Each client $i\in\mathbb{N}$ has its own local dataset $\mathcal{D}_{i}$ with a size of $|\mathcal{D}_{i}|$. Due to limited communication bandwidth, in the $t$-th round, the server typically selects a subset of clients, denoted as $\mathbb{S}^{t}$ with $|\mathbb{S}^{t}|=K<N$. Clients within the subset $\mathbb{S}^{t}$ receive the global model $\bm{w}^{t}$ from the server. They then compute local updates on their respective local datasets, transmitting the resulting local gradients $\nabla f_{j}(\bm{w}^{t})$ back to the server. The server aggregates these gradients and updates the model parameters using
@@
Here, $\eta_{t}$ represents the predefined learning rate, and $\theta^{t}_{j}$ denotes the weight assigned to client $j\in\mathbb{S}^{t}$ in the $t$-th round. The training process lasts for $T$ communication rounds until the global model achieves convergence. Here, the subset $\mathbb{S}^{t}$ and the weight $\theta^{t}_{j}$ with $j\in\mathbb{S}^{t}$ are the variables to be determined.
@@ III-B The Optimization Function
III-B1 Client Selection in a Single Round
In a single round, we identify the subset of clients that best approximates the data distribution of the full set of clients. Specifically, we evaluate the estimation error between the aggregated gradient obtained from the client subset $\mathbb{S}^{t}$ and the aggregated gradient obtained from the full client set $\mathbb{N}$. A small estimation error indicates that the subset of selected clients can effectively represent the data distribution of the full client set. Mathematically, we formulate it as
@@ III-B1 Client Selection in a Single Round
In Eq. (2), when clients are not selected in the $t$-th round, the gradients $\nabla f_{i}(\bm{w}^{t})$ for these clients become unknown, rendering the optimization of $\text{D}(\mathbb{S}^{t})$ intractable. To address this challenge, we obtain the upper bound of $\text{D}(\mathbb{S}^{t})$, denoted as $\text{DUB}(\mathbb{S}^{t})$ in Eq. (3), and minimize this upper bound instead. The proof of Theorem 1 is provided in the supplementary file.
@@
$\text{DUB}(\mathbb{S}^{t})\triangleq\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2},$ (3)
@@
Specifically, following the analysis in [32], we assume a mapping $\xi^{t}:\mathbb{N}\rightarrow\mathbb{S}^{t}$ that assigns each client $i\in\mathbb{N}$ to a client $j\in\mathbb{S}^{t}$, i.e., $\xi^{t}(i)=j$, indicating that client $i$ can be approximately represented by client $j$ in the $t$-th round. For a client $j\in\mathbb{S}^{t}$, let $\mathbb{C}^{t}_{j}=\{i\in\mathbb{N}\,|\,\xi^{t}(i)=j\}$ be the set of clients that can be represented by client $j$. The value $\theta^{t}_{j}=|\mathbb{C}^{t}_{j}|$ is the number of such clients being represented, and is used as the weight of client $j$ in Eq. (1). Based on the mapping $\xi^{t}$, we obtain the upper bound of the estimation error in Eq. (2), as stated in Theorem 1. In the later section, we utilize $\text{DUB}(\mathbb{S}^{t})$ when minimizing the estimation error $\text{D}(\mathbb{S}^{t})$ is required.
To manage the computation of $\text{DUB}(\mathbb{S}^{t})$ in Eq. (3), we introduce the metric $\text{Dist}_{i,j}(t)$ to quantify the distance between two clients $i$ and $j$ in the $t$-th round. This metric is defined as
$\text{Dist}_{i,j}(t)=||\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})||_{2}^{2}.$ (4)
Note that when clients $i,j\in\mathbb{N}$ are not selected in the $t$-th round, their gradients $\nabla f_{i}(\bm{w}^{t})$ and $\nabla f_{j}(\bm{w}^{t})$, and consequently $\text{Dist}_{i,j}(t)$, cannot be calculated. To address this issue, we propose partial updates for $\text{Dist}_{i,j}(t)$. Specifically, when $t=0$, all clients are selected, and we compute $\text{Dist}_{i,j}(t)$ for each pair of clients. Then, for $t\geq 1$, the server updates $\text{Dist}_{i,j}(t)$ using the gradients only from the selected $K$ clients. That is expressed as
@@
Although using $\text{Dist}_{i,j}(t-1)$ to approximate $\text{Dist}_{i,j}(t)$ may introduce biases when either client is not selected, experimental results demonstrate that there are minimal impacts on the model convergence. This can be attributed to two reasons. First, there are limited variations in gradients between successive communication rounds. Moreover, the individual fairness constraint proposed in Section III-B2 ensures that each client has the opportunity of being selected and the corresponding $\text{Dist}_{i,j}(t)$ can be updated. In the following, we use $\text{Dist}_{i,j}(t)$ in Eq. (5) to replace the term $\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2}$ in Eq. (3) to make $\text{DUB}(\mathbb{S}^{t})$ solvable.
III-B2 Client Selection Across Multiple Rounds
In our work, we propose a novel fairness constraint to guide the client selection process across multiple rounds. This constraint asserts that clients with similar data distributions should have similar frequencies of being selected.
To formulate the fairness constraint, we establish metrics to measure the similarity between a pair of clients in terms of their data distributions and the frequencies of being selected, respectively. First, we use $\text{Dist}_{i,j}(t)$ in Eq. (4) to measure the similarity of the data distribution and use Eq. (5) to compute its value. Second, let $x_{i,t}$ represent whether the $i$-th client is selected in the $t$-th round, with $x_{i,t}=1$ if $i\in\mathbb{S}^{t}$ and $x_{i,t}=0$ otherwise. We use

### v2

# Emulating Full Participation: An Effective and Fair Client Selection Strategy for Federated Learning

@@ Abstract
In federated learning, client selection is a critical problem that significantly impacts both model performance and fairness. Prior studies typically treat these two objectives separately, or balance them using simple weighting schemes. However, we observe that commonly used metrics for model performance and fairness often conflict with each other, and a straightforward weighted combination is insufficient to capture their complex interactions. To address this, we first propose two guiding principles that directly tackle the inherent conflict between the two metrics while reinforcing each other. Based on these principles, we formulate the client selection problem as a long-term optimization task, leveraging the Lyapunov function and the submodular nature of the problem to solve it effectively. Experiments show that the proposed method improves both model performance and fairness, guiding the system to converge comparably to full client participation. This improvement can be attributed to the fact that both model performance and fairness benefit from the diversity of the selected clients’ data distributions. Our approach adaptively enhances this diversity by selecting clients based on their data distributions, thereby improving both model performance and fairness.
@@
Federated Learning, Client Selection, Coreset Selection, Individual Fairness, Lyapunov Function
@@ I Introduction
Federated learning (FL) facilitates collaborative model training without the necessity of sharing local data [1, 2] and is widely used in various domains [3, 4]. In FL, model parameters or gradient updates are frequently exchanged between the server and clients, which leads to substantial communication overhead. To address the challenge of limited bandwidth, a common approach is to select a subset of clients for local training [5, 6]. Therefore, a critical challenge is how to select proper and representative clients to participate, which is known as the client selection problem.
In heterogeneous scenarios—where clients hold training data with diverse distributions—the client selection problem becomes particularly critical. Specifically, client selection impacts the FL system in two key aspects. First, client selection influences model performance. Different selection strategies lead to distinct optimization trajectories, and an improper client selection may cause the optimization process to deviate significantly from the optimal path. Second, client selection affects fairness. The global model tends to perform better for frequently selected clients, as their data is better optimized, while producing biased predictions for less frequently selected clients. Since every client participates in the system with the expectation of obtaining accurate predictions for its own data, this selection imbalance may drive underrepresented clients to leave the system. Therefore, it is essential to establish an effective and fair client selection strategy.
Limitation of Prior Works. Prior studies usually treat model performance and fairness as separate objectives and attempt to balance the two using simple weighting schemes [7, 8, 9]. However, model performance and fairness metrics commonly in use are often conflicting, and a simple weighted combination is insufficient to capture their intricate interactions. Specifically, to achieve high model performance, existing client selection methods [10, 11, 12] often prioritize clients with higher training losses, as these clients are considered more challenging to fit their local data. However, such loss-guided methods can lead to biased predictions for clients that are rarely selected. On the other hand, to enhance fairness, existing approaches typically focus on performance fairness [13, 14, 7, 8, 9], which requires the model to deliver similar performance or uniform selection probabilities across all clients. However, this uniform selection approach overlooks the heterogeneous nature of clients’ data distributions, thereby sacrificing model performance. Therefore, a critical challenge lies in reconciling the conflicting goals of model performance and fairness, particularly in heterogeneous scenarios.
Figure 1: Visualization of the client selection results. Clients are represented by blue dots, and selected clients are marked with orange stars. In the non-clustering scenario, clients arranged in a circle are effectively represented by the selected client positioned at the center.
To design an effective and fair client selection strategy, we propose two guiding principles. These principles address the inherent conflict between the two metrics by capturing their interactions. First, from the perspective of model performance, we propose Principle I: The data distribution generated by the selected subset of clients should closely resemble the data distribution of the full client participation. Full client participation is chosen as the standard because models trained with all clients generally yield robust results, serving as a comprehensive benchmark—except in certain extreme cases where some clients possess highly noisy or corrupted data that can skew the model’s performance. Importantly, for clients that are less frequently selected in the loss-guided methods [10, 11, 12], this strategy ensures they are not excluded from selection by approximating the distribution of full client participation.
Furthermore, from the perspective of fairness, we introduce Principle II: clients with similar data distributions should have similar frequencies of being selected. This principle aligns with the concept of Individual Fairness (IF) [15], which posits that instances with similar features in a dataset should yield similar predictions or outcomes. We refer to this principle as the individual fairness constraint. Unlike the uniform sampling strategies used in prior works [8, 9, 13], this constraint specifically focuses on clients with similar data distributions. It is better suited to heterogeneous data scenarios, as it avoids sacrificing model performance by preventing the uniform selection of clients who contribute minimally to model improvement.
The two principles are not independent and mutually reinforce each other. Consider an extreme case where there are two clusters of clients, and the data distributions are similar within each cluster. If only Principle II is applied, it may result in the system selecting clients exclusively from one cluster while neglecting the other. This phenomenon was observed in our experimental results, whereas Principle I helps mitigate this issue. On the other hand, if only Principle I is applied, the system may end up selecting the same client from each cluster repeatedly, whereas Principle II prevents this bias.
In this study, we integrate the two principles and propose an effective and fair client selection strategy for federated learning, called LongFed. We begin by providing a mathematical formulation for both Principle I and Principle II. Then, we model the client selection problem as a long-term optimization function, introducing a tradeoff factor to balance the two principles. To solve this optimization, we simplify it using the Lyapunov optimization from control theory, and propose a fast greedy algorithm based on the submodular nature of the problem. We also theoretically analyze the convergence of the proposed strategy, demonstrating that it converges at a rate of $\mathcal{O}(1/t)$, which is the same as loss-guided selection methods [10, 11, 12, 16].
We evaluate the proposed strategy through extensive experiments, with results demonstrating that our method enhances both model performance and fairness. Regarding model performance, the strategy effectively guides the system to converge along a trajectory similar to that of full client participation, outperforming prior methods in most cases. Additionally, the strategy achieves strong fairness, evidenced by a low standard deviation in the selection frequencies of clients with similar data distributions. This improvement can be attributed to the fact that both model performance and fairness benefit from the diversity of the selected clients’ data distributions, which is promoted by our two proposed principles. Unlike existing fair federated learning approaches [8, 9, 13] that treat all clients equally, our method adaptively enhances this diversity based on clients’ data distributions, making it better suited for heterogeneous environments. Additionally, our method introduces only a marginal time increase (less than 0.4 ms) compared to existing approaches.
We also provide visualization results to illustrate the effectiveness of our method. An example is shown in Fig. 1. When clients exhibit clear clustering patterns, the proposed strategy selects one client from each cluster and chooses different clients across multiple rounds. This selection process aligns with the clustered federated learning [17, 18]. More importantly, in scenarios where the clustering pattern is not evident—which is more common in practical cases—the proposed strategy selects clients that cover a majority of client population. This diversity not only enhances model performance but also improves fairness.
Our contributions are as follows.
  • We identify the inherent conflict between model performance and fairness in the client selection problem and propose a strategy that leads to improvements in both.
@@
  • We address a previously underexplored issue: the frequency of selecting clients with similar data distributions. We introduce an individual fairness criterion to mathematically formulate and effectively resolve this issue.
@@ II-B Fairness in Federated Learning
Performance fairness refers that the model should produce similar performance across all clients, aligning most closely with our objectives. Various approaches have emerged to achieve performance fairness. For example, in [7], fairness is defined by assigning uniform weight to each client and introduced as an additional constraint in the optimization function. Besides, [14] identifies a trade-off between model robustness and fairness, and proposes a personalized framework to inherently achieve both fairness and robustness benefits. Moreover, several works study the fairness issue by considering that the probability of being selected is similar for all clients [8, 9, 29]. However, as introduced in Section I, the uniform selection method disregards the heterogeneous data distribution among clients.
In summary, existing client selection methods only consider the model performance and do not address the fairness. Although some fair federated learning methods have been proposed, the uniform selection constraint adopted in their methods could disregard the heterogeneous data distribution among clients. Therefore, it remains a critical problem for client selection to simultaneously address model performance and fairness.
@@ III The Proposed Optimization Function
In this section, we begin with an introduction to the federated learning system. Next, we provide a mathematical formulation for both Principle I and Principle II, and model the client selection problem as a long-term optimization function. Last, we apply Lyapunov optimization to simplify the optimization function.
@@ III-A Preliminary of Federated Learning
In our work, we consider that the federated learning system consists of a central server and a set of clients denoted as =$\mathbb{N}=\{1,\cdots,N\}$. Each client $i\in\mathbb{N}$ has its own local dataset $\mathcal{D}_{i}$ with a size of $|\mathcal{D}_{i}|$. In the $t$-th round, the server selects a subset of clients, denoted as $\mathbb{S}^{t}$ with $|\mathbb{S}^{t}|=K<N$. Clients within the subset $\mathbb{S}^{t}$ receive the global model $\bm{w}^{t}$ from the server. They then compute local updates on their respective local datasets, transmitting the local gradients $\nabla f_{j}(\bm{w}^{t})$ back to the server. The server aggregates these gradients and updates the model using
@@
Here, $\eta_{t}$ represents the predefined learning rate, and $\theta^{t}_{j}$ denotes the weight assigned to client $j\in\mathbb{S}^{t}$ in the $t$-th round. The training process lasts for $T$ rounds until the global model achieves convergence. The subset $\mathbb{S}^{t}$ and the weights $\theta^{t}_{j}$ with $j\in\mathbb{S}^{t}$ are the variables to be determined.
@@ III-B The Optimization Function
III-B1 Formulation of Principle I
In our work, we employ the Principle I to guide the client selection in a single round. Specifically, we evaluate the estimation error between the aggregated gradient obtained from the client subset $\mathbb{S}^{t}$ and the aggregated gradient obtained from the full client set $\mathbb{N}$. A small estimation error indicates that the subset of selected clients can effectively represent the data distribution of the full client set. Mathematically, we formulate it as
@@ III-B1 Formulation of Principle I
In Eq. (2), when clients are not selected in the $t$-th round, their gradients $\nabla f_{i}(\bm{w}^{t})$ become unknown. To address this challenge, following the analysis in [30], we assume a mapping $\xi^{t}:\mathbb{N}\rightarrow\mathbb{S}^{t}$ that assigns each client $i\in\mathbb{N}$ to a client $j\in\mathbb{S}^{t}$, i.e., $\xi^{t}(i)=j$, indicating that client $i$ can be approximately represented by client $j$ in the $t$-th round. For a client $j\in\mathbb{S}^{t}$, let $\mathbb{C}^{t}_{j}=\{i\in\mathbb{N}\,|\,\xi^{t}(i)=j\}$ be the set of clients that can be represented by client $j$. The value $\theta^{t}_{j}=|\mathbb{C}^{t}_{j}|$ is the number of such clients being represented, and is used as the weight of client $j$ in Eq. (1). Based on the mapping $\xi^{t}$, we obtain the upper bound of the estimation error in Eq. (2), as stated in Theorem 1. In the later section, we utilize $\text{DUB}(\mathbb{S}^{t})$ when minimizing the estimation error $\text{D}(\mathbb{S}^{t})$ is required. The proof of Theorem 1 is provided in supplementary file.
@@
$\text{Dist}_{i,j}(t)=||\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})||_{2}^{2},$ (3)
and
$\text{DUB}(\mathbb{S}^{t})$ $\triangleq\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\text{Dist}_{i,j}(t),$ (4)
@@
In our work, we use partial updates to compute $\text{Dist}_{i,j}(t)$. Specifically, when $t=0$, all clients are selected, and we compute $\text{Dist}_{i,j}(t)$ for each pair of clients. Then, for $t\geq 1$, the server updates $\text{Dist}_{i,j}(t)$ using the gradients only from the selected $K$ clients. That is,
@@
Although using $\text{Dist}_{i,j}(t-1)$ to approximate $\text{Dist}_{i,j}(t)$ may introduce biases, experimental results demonstrate that there are minimal impacts on the model convergence. This is because there is limited gradient variations between successive communication rounds, and the proposed individual fairness constraint ensures that all clients have the opportunity of being selected and the corresponding $\text{Dist}_{i,j}(t)$ can be updated.
III-B2 Formulation of Principle II
As mentioned in Section I, we propose the individual fairness constraint (Principle II) to guide the client selection across multiple rounds. It asserts that clients with similar data distributions should have similar frequencies of being selected.
First, we use $\text{Dist}_{i,j}(t)$ in Eq. (3) to measure the similarity of the data distribution. Second, let $x_{i,t}$ represent whether the $i$-th client is selected in the $t$-th round, with $x_{i,t}=1$ if $i\in\mathbb{S}^{t}$ and $x_{i,t}=0$ otherwise. We use